# Scalable Feature Learning on Huge Knowledge Graphs for Downstream Machine Learning

### v1

@@ Abstract
Many machine learning tasks can benefit from external knowledge. Large knowledge graphs store such knowledge, and embedding methods can be used to distill it into ready-to-use vector representations for downstream applications. For this purpose, current models have however two limitations: they are primarily optimized for link prediction, via local contrastive learning, and they struggle to scale to the largest graphs due to GPU memory limits. To address these, we introduce SEPAL: a Scalable Embedding Propagation ALgorithm for large knowledge graphs designed to produce high-quality embeddings for downstream tasks at scale. The key idea of SEPAL is to enforce global embedding alignment by optimizing embeddings only on a small core of entities, and then propagating them to the rest of the graph via message passing. We evaluate SEPAL on 7 large-scale knowledge graphs and 46 downstream machine learning tasks. Our results show that SEPAL significantly outperforms previous methods on downstream tasks. In addition, SEPAL scales up its base embedding model, enabling fitting huge knowledge graphs on commodity hardware.
@@ 1 Introduction: embedding knowledge for downstream tasks
Bringing general knowledge to a machine-learning task revives an old promise of making it easier via this knowledge [Lenat and Feigenbaum, 2000]. Data science is often about entities of the world –persons, places, organizations– that are increasingly well characterized in general-purpose knowledge graphs. These graphs carry rich information, including numerical attributes and relationships between entities, and can be connected to string values in tabular data through entity linking techniques [Mendes et al., 2011, Foppiano and Romary, 2020, Delpeuch, 2019]. However, integrating the corresponding data into tabular machine learning often raises the thorny challenge of assembling features from a relational graph [Kanter and Veeramachaneni, 2015, Cappuzzo et al., 2025, Robinson et al., 2025]. Graph embedding methods offer a solution by distilling the graph information into node features readily usable by any downstream tabular learner [Grover and Leskovec, 2016, Cvetkov-Iliev et al., 2023, Ruiz et al., 2024].
@@
The rapid growth of general-purpose knowledge graphs brings the exciting prospect of a very general feature enrichment. Indeed, the richer the knowledge graph, the more it can provide value to the downstream analysis [Ruiz et al., 2024]. ConceptNet pioneered the distribution of general-knowledge embeddings, building on a graph of 8 million entities [Speer et al., 2017]. As of 2025, Wikidata [Vrandečić and Krötzsch, 2014] describes 115M entities, and YAGO4 gives a curated view on 67M entities [Pellissier Tanon et al., 2020]. Moreover, these knowledge graphs are rapidly expanding, with Wikidata alone gaining around 15M entities yearly [Wikimedia, ].
In parallel, the sophistication of knowledge-graph embedding (KGE) models is increasing [Bordes et al., 2013, Yang et al., 2015, Balazevic et al., 2019b, …], capturing better the relational aspect of the data, important for downstream tasks [Cvetkov-Iliev et al., 2023]. However, much of the KGE literature remains focused on link prediction, a task shown to correlate poorly with the utility of the embeddings on real-world downstream tasks [Ruffinelli and Gemulla, 2024]. One reason may be that, for link prediction, models typically optimize for local contrasts, resulting in embeddings that are not calibrated [Tabacof and Costabello, 2020].
Another challenge is that embedding models do not scale easily. There can be a disconnect between the growth of knowledge graphs, and the fact that increasingly complex embedding models tend to be less tractable, and are typically demonstrated on comparatively small graphs, subsets of real-world graphs such as FB15k (15k entities from Freebase) or WN18 (40k entities from WordNet), three orders of magnitude smaller than modern general knowledge graphs or industrial knowledge graphs [Sullivan, 2020]. The typical answer to this challenge is distributed computation across GPUs or machines, explored by PyTorch-BigGraph [Lerer et al., 2019] and many more [Zheng et al., 2020, Mohoney et al., 2021, Zhu et al., 2019, Dong et al., 2022, Ren et al., 2022, Zheng et al., 2024, …]. This comes with sizeable computational costs, as the graph is fit piece by piece, increasing the data movement across devices for entities involved in several edge partitions. Moreover, the hardware requirement is a challenge for non-profits such as the Wikimedia Foundation, in charge of Wikidata, given that the embeddings should be recomputed regularly to incorporate newly added entities.
Here we present a KGE approach to 1) capture better some global structure of the graph, 2) scale to huge graphs with little computational resources. This approach applies as a wrapper to many KGE methods. It bridges the gap between advances in embedding approaches and the goal to create large and reusable general-knowledge embeddings for downstream applications. For this, we leverage a fundamental structure in knowledge graphs: a small set of core entities comes with much more information than the others. Our method, SEPAL (Scalable Embedding Propagation ALgorithm), scales up any embedding model that, at triple-level, models the tail embedding as a relation-specific transformation of the head embedding. Two technical contributions are key to SEPAL’s scalability. 1) Time-wise: we show that message passing, rather than classic optimization, can be used to propagate embeddings of a small core subset of entities and give good embeddings for the full graph. The challenge here is to maintain the relational geometry. 2) Memory-wise: we devise an algorithm called BLOCS to break down a huge knowledge graph into overlapping subsets that fit in GPU memory. Here, the challenge lies in the scale-free and connectivity properties of a large knowledge graph: some nodes are connected to a significant fraction of the graph, while others are very hard to reach.
We start by reviewing related work. Then, section 3 describes our contributed method and section 4 gives a theoretical analysis. In section 5 we evaluate SEPAL’s performance on knowledge graphs of increasing size between YAGO3 [2.6M entities, Mahdisoltani et al., 2014] and WikiKG90Mv2 [91M entities, Hu et al., 2020]; we study the use of the embeddings for feature enrichment on 46 downstream machine learning tasks, showing that SEPAL makes embedding methods more tractable while generating better embeddings for downstream tasks. Finally, section 6 discusses the contributions and limitations of SEPAL. Empirical findings show that:
@@
    SEPAL’s propagation of core embeddings to outer entities with message passing improves the downstream performance compared to classic optimization;
@@
    this approach can be scaled to very large knowledge graphs on modest hardware. We embed WikiKG90Mv2 –91M entities, 601M triples– with a single 32GB V100 GPU.
2 Related work: embedding and scalability
@@
2.1 Embedding knowledge graphs
Here, we provide an overview of approaches to generate low-dimensional (typically $d=100$ or $200$) vector representations for the entities, that can be used in downstream applications. These include both graph-embedding techniques, that leave aside the relations, and knowledge-graph embedding (KGE) methods accounting for relations.
@@ 2.1 Embedding knowledge graphs
A first simple strategy to get very cost-effective vector representations is to compute random projections. This avoids relying on –potentially costly– optimization, and provides embeddings preserving the pairwise distances to within an epsilon [Dasgupta and Gupta, 2003]. FastRP [Chen et al., 2019a] proposes a scalable approach, with a few well-chosen very sparse random projections of the normalized adjacency matrix and its powers.
Another family of methods performs explicit matrix factorizations on matrices derived from the adjacency matrix, for instance GraREP [Cao et al., 2015] or NetMF [Qiu et al., 2018]. These methods output close embedding vectors for nodes with similar neighborhoods.
Similarly, Skip-Gram Negative Sampling (SGNS), behind word2vec [Mikolov et al., 2013a, b], performs an implicit factorization [Levy and Goldberg, 2014]. It has been adapted to graphs: DeepWalk [Perozzi et al., 2014] and node2vec [Grover and Leskovec, 2016] use random walks on the graph to generate “sentences” fed to word2vec. LINE [Tang et al., 2015] explores a similar strategy varying edge sampling.
@@
RDF2vec [Ristoski and Paulheim, 2016] adapts SGNS to multi-relational graphs by simply adding the relations to the generated sentences.
@@
where $\phi$ is a model-specific relational operator, and $\mathrm{sim}$ a similarity function. The embeddings are optimized by gradient descent to maximize the score of positive triples, and minimize that of negative ones. These models strive to align, for positive triples, the tail embedding $\boldsymbol{\theta}_{t}$ with the “relationally” transformed head embedding $\phi(\boldsymbol{\theta}_{h},\boldsymbol{w}_{r})$. The challenge is to design a clever $\phi$ operator to model complex patterns in the data –hierarchies, compositions, symmetries… Indeed some relations are one-to-one (people only have one biological mother), well represented by a translation [Bordes et al., 2013], while others are many-to-one (for instance many person were BornIn Paris), calling for $\phi$ to contract distances [Wang et al., 2017]. Many models explore different parametrizations, among which MuRE [Balazevic et al., 2019b], RotatE [Sun et al., 2019], or QuatE [Zhang et al., 2019] have good performance [Ali et al., 2021a]. This framework also includes models like DistMult [Yang et al., 2015], ComplEx [Trouillon et al., 2016], or TuckER [Balazevic et al., 2019a], that implicitly perform tensor factorizations.
Table 1: Expression of $\phi$ in some embedding models. $\odot$ denotes the Hadamard product, $\otimes$ the Hamilton product, and ×$\times_{i}$ the tensor product along mode $i$. The models we list here are all compatible with our proposed SEPAL approach.
Model Relational operator $\phi$
TransE [Bordes et al., 2013] $\boldsymbol{\theta}_{h}+\boldsymbol{w}_{r}$
MuRE [Balazevic et al., 2019b] $\boldsymbol{\theta}_{h}\odot\boldsymbol{\rho}_{r}-\boldsymbol{w}_{r}$
RotatE [Sun et al., 2019] $\boldsymbol{\theta}_{h}\odot\boldsymbol{w}_{r}$
QuatE [Zhang et al., 2019] $\boldsymbol{\theta}_{h}\otimes\boldsymbol{w}_{r}$
DistMult [Yang et al., 2015] $\boldsymbol{\theta}_{h}\odot\boldsymbol{w}_{r}$
ComplEx [Trouillon et al., 2016] $\boldsymbol{\theta}_{h}\odot\boldsymbol{w}_{r}$
TuckER [Balazevic et al., 2019a] $\boldsymbol{\mathcal{W}}\times_{1}\boldsymbol{\theta}_{h}\times_{2}\boldsymbol{w}_{r}$
@@
To smooth computed embeddings, CompGCN [Vashishth et al., 2020] introduces the idea of propagating knowledge-graph embeddings using the relational operator $\phi$, but couples it with learnable weights and a non-linearity. REP [Wang et al., 2022] simplifies this framework by removing weight matrices and non-linearities. Rossi et al. [2022] also uses Feature Propagation, but to impute missing node features in graphs.
@@ 2.2 Techniques for scaling graph algorithms
Scaling up graph computations, for graph embedding or more generally, often relies on breaking down graphs in subgraphs. Section E.1 presents corresponding prior work.
@@
Other forms of data reduction can help to scale graph algorithms (e.g. based on message passing). Algorithms may subsample neighborhoods, as GraphSAGE [Hamilton et al., 2017] that selects a fixed number of neighbors for each node on each layer, or MariusGNN [Waleffe et al., 2023] that uses an optimized data structure for neighbor sampling and GNN aggregation. Cluster-GCN [Chiang et al., 2019] restricts the neighborhood search within clusters, obtained by classic clustering algorithms, to improve computational efficiency. GraphSAINT [Zeng et al., 2020] samples overlapping subgraphs through random walks, for supervised GNN training via node classification, optimizing GNN weights without processing the full graph.
@@
Multi-level approaches, such as HARP [Chen et al., 2018], GraphZoom [Deng et al., 2020] or MILE [Liang et al., 2021], coarsen the graph, compute embeddings on the smaller graph, and project them back to the original graph.
@@ 2.3 Scaling knowledge-graph embedding
Many approaches scale triple-level stochastic solvers by distributing training across multiple workers, starting from the seminal PyTorch-BigGraph (PBG) [Lerer et al., 2019] that splits the triples into buckets based on the partitioning of the entities. The challenge is then to limit overheads and communication costs coming from 1) the additional data movement incurred by embeddings of entities occurring in several buckets 2) the synchronization of global trainable parameters such as the relation embeddings. For this, DGL-KE [Zheng et al., 2020] reduces data movement by using sparse relation embeddings and METIS graph partitioning [Karypis and Kumar, 1997] to distribute the triples across workers. HET-KG [Dong et al., 2022] further optimizes distributed training by preserving a copy of the most frequently used embeddings on each worker, to reduce communication costs. These “hot-embeddings” are periodically synchronized to minimize inconsistency. SMORE [Ren et al., 2022] leverages asynchronous scheduling to overlap CPU-based data sampling, with GPU-based embedding computations. Algorithmically, it contributes a rejection sampling strategy to generate the negatives at low cost. GraphVite [Zhu et al., 2019] accelerates SGNS for graph embedding by both parallelizing random walk sampling on multiple CPUs, and negative sampling on multiple GPUs. Marius [Mohoney et al., 2021] reduces synchronization overheads by opting for asynchronous training of entity embeddings with bounded staleness, and minimizes IO with partition caching and buffer-aware data ordering. GE2 [Zheng et al., 2024] improves data swap between CPU and multiple GPUs. Finally, the LibKGE [Broscheit et al., 2020] Python library also supports parallel training and includes GraSH [Kochsiek et al., 2022], an efficient hyperparameter optimization algorithm for large-scale KGE models.
@@
Other approaches reduce GPU memory pressure by limiting the number of parameters. NodePiece [Galkin et al., 2022] and EARL [Chen et al., 2023] embed a subset of entities and train an encoder to compute the embeddings of the other entities. However, their tokenization step is costly, and they have not been demonstrated on graphs larger than 2.5M nodes.
3 SEPAL: expanding from a core subgraph

### v2

@@ Abstract
Many machine learning tasks can benefit from external knowledge. Large knowledge graphs store such knowledge, and embedding methods can be used to distill it into ready-to-use vector representations for downstream applications. For this purpose, current models have however two limitations: they are primarily optimized for link prediction, via local contrastive learning, and their application to the largest graphs requires significant engineering effort due to GPU memory limits. To address these, we introduce SEPAL: a Scalable Embedding Propagation ALgorithm for large knowledge graphs designed to produce high-quality embeddings for downstream tasks at scale. The key idea of SEPAL is to ensure global embedding consistency by optimizing embeddings only on a small core of entities, and then propagating them to the rest of the graph with message passing. We evaluate SEPAL on 7 large-scale knowledge graphs and 46 downstream machine learning tasks. Our results show that SEPAL significantly outperforms previous methods on downstream tasks. In addition, SEPAL scales up its base embedding model, enabling fitting huge knowledge graphs on commodity hardware. Our code is available at: https://github.com/flefebv/sepal.git.
@@ 1 Introduction: embedding knowledge for downstream tasks
Bringing general knowledge to a machine-learning task revives an old promise of making it easier via this knowledge (lenat2000thresholds). Indeed, data science is often about entities of the world –persons, places, organizations– that are well characterized in general-purpose knowledge graphs. These graphs carry rich information, including numerical attributes and relationships between entities, and can be connected to string values in tabular data through entity linking techniques (mendes2011dbpedia; foppiano2020entity; delpeuch2019opentapioca). A thorny challenge, however, is to transform this relational information into features for downstream tabular machine learning (kanter2015deep; cappuzzo2024retrieve; robinson2025relbench). To that end, a scalable solution is offered by graph embedding methods that distill the graph information into node features readily usable by any downstream tabular learner (grover2016node2vec; cvetkov2023relational; ruiz2024high).
Knowledge graphs as general knowledge sources
@@
The rapid growth of general-purpose knowledge graphs brings the exciting prospect of a very general feature enrichment. Indeed, a richer knowledge graph provides more comprehensive coverage and context, thereby bringing greater value to the downstream analysis (ruiz2024high). ConceptNet pioneered the distribution of general-knowledge embeddings, building on a graph of 8 million entities (speer2017conceptnet). Since then, knowledge graphs have continued to expand rapidly. For instance, as of 2025, Wikidata (vrandevcic2014wikidata) describes 115M entities and gains around 15M yearly (wikidatagrowth), and YAGO4 (pellissier2020yago) gives a curated view on 67M entities.
Optimizing embeddings for the right task
In parallel, the sophistication of knowledge-graph embedding (KGE) models is increasing (bordes2013translating; yang2015embedding; balazevic2019multi), capturing better the relational aspect of the data, important for downstream tasks (cvetkov2023relational). However, most of the KGE literature prioritizes link prediction as the primary benchmark, despite recent findings showing that strong performance on this task does not correlate with improved performance on downstream predictive tasks (ruffinelli2024beyond). One reason may be that, for link prediction, models typically optimize for local contrasts, resulting in embeddings that are not calibrated (Tabacof2020Probability; arakelyan2023adapting). While prior work has explored multi-hop reasoning to capture more complex graph patterns (hamilton2018embedding; ren2020beta), the standard evaluation paradigm for KGEs still revolves around internal tasks, rather than how the learned embeddings can transfer knowledge to practical machine-learning tasks beyond the knowledge graph itself.
The importance of scalability
To leverage the full potential of very large knowledge graphs, embedding methods must be highly scalable. While many methods have been proposed to scale KGE models, doing so is not trivial. Sophisticated KGE models are typically demonstrated on small datasets like FB15k (15k entities) or WN18 (40k entities), which are orders of magnitude smaller than modern general-purpose or industrial knowledge graphs (sullivan2020reintroduction). The common solution to this scalability challenge is either distributed computation across multiple GPUs or machines (lerer2019pytorch; zhu2019graphvite; zheng2020dgl; dong2022het; zheng2024ge2), or leveraging the full memory hierarchy (disk, CPU, and GPU) on a single machine (mohoney2021marius; ren2022smore). These approaches require significant engineering effort to manage data partitioning, optimize data movement, and minimize synchronization overheads.
Contributions
In this paper, we aim to bridge the gap between advances in embedding methods and the goal to create large and reusable general-knowledge embeddings for downstream applications. We introduce SEPAL, a scalable algorithm that applies as a wrapper to many embedding models. Our contributions are:
@@
    We propose a new embedding optimization strategy that enforces global consistency. Instead of optimizing all embeddings with local contrastive learning, SEPAL first processes a small but dense core of the graph, to learn relation and core-entity embeddings. It then propagates these embeddings to the remaining entities using relation-aware message passing. The absence of negative sampling at this propagation stage accelerates the embedding computation and makes them better suited for downstream tasks.
@@
    We provide a theoretical analysis showing that SEPAL’s propagation step, combined with DistMult, implicitly maximizes the alignment of embeddings within positive triples.
  3. 3.
    We introduce BLOCS, a scalable graph-splitting algorithm that partitions huge, scale-free graphs into manageable, overlapping subgraphs. This enables fitting the embedding process on a single GPU, avoiding the engineering complexity of distributed systems. Here, the challenge lies in the scale-free and connectivity properties of a large knowledge graph: some nodes are connected to a significant fraction of the graph, while others are hard to reach.
  4. 4.
    We conduct an extensive empirical study on 7 large knowledge graphs and 46 downstream tasks. Results show that SEPAL significantly outperforms standard methods on downstream tasks and is generally faster than existing large-scale systems. Moreover, it scales to ultra-large graphs with little computational resources: we embed WikiKG90Mv2 –91M entities, 601M triples– with a single 32GB V100 GPU. Our experimental results also highlight that using such large knowledge graphs is beneficial for downstream tasks in real-world feature enrichment scenarios.
We start by reviewing related work in section 2. Then, section 3 describes our contributed method and section 4 gives a theoretical analysis. In section 5 we evaluate SEPAL’s performance on knowledge graphs of increasing size between YAGO3 (2.6M entities, mahdisoltani2014yago3) and WikiKG90Mv2 (91M entities, hu2020open); we study the use of the embeddings for feature enrichment on 46 downstream machine learning tasks, showing that SEPAL makes embedding methods more tractable while generating better embeddings for downstream tasks. Finally, section 6 discusses the contributions and limitations of SEPAL.
2 Related work: embedding optimization and scalability
@@
2.1 Optimizing knowledge-graph embeddings
Here, we provide an overview of approaches to generate low-dimensional (typically $d=100$) vector representations for the entities, that can be used in downstream applications. These include both graph-embedding techniques, that leave aside the relations, and KGE methods accounting for relations.
@@ 2.1 Optimizing knowledge-graph embeddings
A first simple strategy to get very cost-effective vector representations is to compute random projections. This avoids relying on –potentially costly– optimization, and provides embeddings preserving the pairwise distances to within an epsilon (dasgupta2003elementary). FastRP (chen2019fast) proposes a scalable approach, with a few well-chosen very sparse random projections of the normalized adjacency matrix and its powers.
Another family of methods performs explicit matrix factorizations on matrices derived from the adjacency matrix, for instance GraREP (cao2015grarep) or NetMF (qiu2018network). These methods output close embedding vectors for nodes with similar neighborhoods.
Similarly, Skip-Gram Negative Sampling (SGNS), behind word2vec (mikolov2013efficient; mikolov2013distributed), performs an implicit factorization (levy2014neural). It has been adapted to graphs: DeepWalk (perozzi2014deepwalk) and node2vec (grover2016node2vec) use random walks on the graph to generate “sentences" fed to word2vec. LINE (tang2015line) explores a similar strategy varying edge sampling. Here, the loss function is typically a binary logistic regression objective:
$\mathcal{L}_{\text{SGNS}}=-\log\sigma(\boldsymbol{\theta}_{w_{c}}^{\top}\boldsymbol{\theta}_{w_{t}})-\sum_{i=1}^{p}\log\left(1-\sigma(\boldsymbol{\theta}_{w_{i}}^{\top}\boldsymbol{\theta}_{w_{t}})\right)$ (1)
with $\boldsymbol{\theta}_{w_{t}}$ and $\boldsymbol{\theta}_{w_{c}}$ the embeddings of the target and context nodes (or words), $w_{i}$ the $i$-th negative sample drawn from a noise distribution, $p$ the number of negative samples, and $\sigma$ the sigmoid function.
Table 1: Expression of $\phi$ in some embedding models. $\odot$ denotes the Hadamard product, $\otimes$ the Hamilton product, and $\times_{i}$ the tensor product along mode ii. The models we list here are all compatible with our proposed SEPAL approach.
Model Relational operator $\phi$
TransE (bordes2013translating) $\boldsymbol{\theta}_{h}+\boldsymbol{w}_{r}$
MuRE (balazevic2019multi) $\boldsymbol{\theta}_{h}\odot\boldsymbol{\rho}_{r}-\boldsymbol{w}_{r}$
RotatE (sun2019rotate) $\boldsymbol{\theta}_{h}\odot\boldsymbol{w}_{r}$
QuatE (zhang2019quaternion) $\boldsymbol{\theta}_{h}\otimes\boldsymbol{w}_{r}$
DistMult (yang2015embedding) $\boldsymbol{\theta}_{h}\odot\boldsymbol{w}_{r}$
ComplEx (trouillon2016complex) $\boldsymbol{\theta}_{h}\odot\boldsymbol{w}_{r}$
TuckER (balavzevic2019tucker) $\boldsymbol{\mathcal{W}}\times_{1}\boldsymbol{\theta}_{h}\times_{2}\boldsymbol{w}_{r}$
@@
RDF2vec (ristoski2016rdf2vec) adapts SGNS to multi-relational graphs by simply adding the relations to the generated sentences.
@@
where $\phi$ is a model-specific relational operator, and $\mathrm{sim}$ a similarity function. The embeddings are optimized by gradient descent to maximize the score of positive triples, and minimize that of negative ones. A possible loss function is the binary cross-entropy loss (ali2021bringing)
$\mathcal{L}_{\text{BCE}}=-\log\sigma(f(h,r,t))-\sum_{i=1}^{p}\log\big(1-\sigma(f(h_{i}^{\prime},r,t_{i}^{\prime}))\big)$ (3)
which boils down to SGNS for $f(h,r,t)=\boldsymbol{\theta}_{h}^{\top}\boldsymbol{\theta}_{t}$. These models strive to align, for positive triples, the tail embedding $\boldsymbol{\theta}_{t}$ with the “relationally” transformed head embedding $\phi(\boldsymbol{\theta}_{h},\boldsymbol{w}_{r})$. The challenge is to design a clever $\phi$ operator to model complex patterns in the data, like hierarchies, compositions, or symmetries. Indeed some relations are one-to-one (people only have one biological mother), well represented by a translation (bordes2013translating), while others are many-to-one (for instance many person were BornIn Paris), calling for $\phi$ to contract distances (wang2017knowledge). Many models explore different parametrizations, among which MuRE (balazevic2019multi), RotatE (sun2019rotate), or QuatE (zhang2019quaternion) have good performance (ali2021bringing). This framework also includes models like DistMult (yang2015embedding), ComplEx (trouillon2016complex), or TuckER (balavzevic2019tucker), that implicitly perform tensor factorizations.
@@
To smooth computed embeddings, CompGCN (vashishth2020compositionbased) introduces the idea of propagating knowledge-graph embeddings using the relational operator $\phi$, but couples it with learnable weights and a non-linearity. REP (wang2022rep) simplifies this framework by removing weight matrices and non-linearities. rossi2022unreasonable also use Feature Propagation, but to impute missing node features in graphs. albooyeh2020out incorporate propagation within the standard link prediction pipeline, with negative sampling and gradient descent on standard KGE loss functions.
@@ 2.2 Techniques for scaling graph algorithms
Scaling up graph computations, for graph embedding or more generally, often relies on breaking down graphs in subgraphs. Section˜E.1 presents corresponding prior work.
@@
Other forms of data reduction can help to scale graph algorithms (e.g. based on message passing). Algorithms may subsample neighborhoods, as GraphSAGE (hamilton2017inductive) that selects a fixed number of neighbors for each node on each layer, or MariusGNN (waleffe2023mariusgnn) that uses an optimized data structure for neighbor sampling and GNN aggregation. Cluster-GCN (chiang2019cluster) restricts the neighborhood search within clusters, obtained by classic clustering algorithms, to improve computational efficiency. GraphSAINT (zeng2020graphsaint) samples overlapping subgraphs through random walks, for supervised GNN training via node classification, optimizing GNN weights without processing the full graph.
@@
Multi-level approaches, such as HARP (chen2018harp), GraphZoom (deng2019graphzoom) or MILE (liang2021mile), coarsen the graph, compute embeddings on the smaller graph, and project them back to the original graph.
@@ 2.3 Scaling knowledge-graph embedding
Many approaches scale triple-level stochastic solvers by distributing training across multiple workers, starting from the seminal PyTorch-BigGraph (PBG) (lerer2019pytorch) that splits the triples into buckets based on the partitioning of the entities. The challenge is then to limit overheads and communication costs coming from 1) the additional data movement incurred by embeddings of entities occurring in several buckets 2) the synchronization of global trainable parameters such as the relation embeddings. For this, DGL-KE (zheng2020dgl) reduces data movement by using sparse relation embeddings and METIS graph partitioning (karypis1997metis) to distribute the triples across workers. HET-KG (dong2022het) further optimizes distributed training by preserving a copy of the most frequently used embeddings on each worker, to reduce communication costs. These “hot-embeddings" are periodically synchronized to minimize inconsistency. SMORE (ren2022smore) leverages asynchronous scheduling to overlap CPU-based data sampling, with GPU-based embedding computations. Algorithmically, it contributes a rejection sampling strategy to generate the negatives at low cost. GraphVite (zhu2019graphvite) accelerates SGNS for graph embedding by both parallelizing random walk sampling on multiple CPUs, and negative sampling on multiple GPUs. Marius (mohoney2021marius) reduces synchronization overheads by opting for asynchronous training of entity embeddings with bounded staleness, and minimizes IO with partition caching and buffer-aware data ordering. GE2 (zheng2024ge2) improves data swap between CPU and multiple GPUs. Finally, the LibKGE (libkge) Python library also supports parallel training and includes GraSH (kochsiek2022start), an efficient hyperparameter optimization algorithm for large-scale KGE models.
@@
Other approaches reduce GPU memory pressure by limiting the number of parameters. NodePiece (galkin2021nodepiece) and EARL (chen2023entity) embed a subset of entities and train an encoder to compute the embeddings of the other entities. However, their tokenization step is costly, and they have not been demonstrated on graphs larger than 2.5M nodes.
3 SEPAL: revisiting knowledge-graph embedding optimization